# ref.bib **footnotetext: Equal contribution

### v1

@@ Abstract
The recent developments in Large Multi-modal Video Models (Video-LMMs) have significantly enhanced our ability to interpret and analyze video data. Despite their impressive capabilities, current Video-LMMs have not been evaluated for anomaly detection tasks, which is critical to their deployment in practical scenarios e.g., towards identifying deepfakes, manipulated video content, traffic accidents and crimes. In this paper, we introduce VANE-Bench, a benchmark designed to assess the proficiency of Video-LMMs in detecting and localizing anomalies and inconsistencies in videos. Our dataset comprises an array of videos synthetically generated using existing state-of-the-art text-to-video generation models, encompassing a variety of subtle anomalies and inconsistencies grouped into five categories: unnatural transformations, unnatural appearance, pass-through, disappearance and sudden appearance. Additionally, our benchmark features real-world samples from existing anomaly detection datasets, focusing on crime-related irregularities, atypical pedestrian behavior, and unusual events. The task is structured as a visual question-answering challenge to gauge the models’ ability to accurately detect and localize the anomalies within the videos. We evaluate nine existing Video-LMMs, both open and closed sources, on this benchmarking task and find that most of the models encounter difficulties in effectively identifying the subtle anomalies. In conclusion, our research offers significant insights into the current capabilities of Video-LMMs in the realm of anomaly detection, highlighting the importance of our work in evaluating and improving these models for real-world applications. Our code and data is available at https://hananshafi.github.io/vane-benchmark/.
@@ 1 Introduction
Figure 1: Samples showing the AI-Generated video category of VANE-Bench. We collect these synthetic videos from SORA [sora], Open-Sora [open-sora], Runway Gen2 [gen2-runway], ModelScopeT2V [wang2023modelscope], and VideoLCM [wang2023videolcm]. The correct option in each question is highlighted in bold. Note that many of these anomalies are extremely subtle and difficult for humans to detect since the changes happen in rapid succession, with the entire video played in under a second. Anomalies identified with red bounding boxes for clarity. Note that our actual dataset does not contain bounding box overlays.
@@
Large Language Models (LLMs) like ChatGPT have ushered in a new era of real-world AI applications in varied and diverse sectors like manufacturing, legal services, space exploration, transportation, retail, healthcare, education, and technology [abdullah2022chatgpt, Marr2023]. Further, the current trend in development of these LLMs have been to introduce multi-modal capabilities like vision, and audio to these models along with text [geminiteam2024gemini, gpt4-o]. This motivates us to ask the question whether the current Large Multi-modal Models (LMMs) are capable and accurate in tackling the problem statement of Video Anomaly Detection (VAD) which has immense practical applications in factories, autonomous driving, crime warning, and traffic management [vad-survey].
@@
Further, we have recently observed superior visual quality of various AI-generated videos due to the improvements in the underlying algorithms, which are based on diffusion models, and transformers [sora, open-sora, peebles2023scalable]. The current state-of-the-art (SOTA) AI text-to-video model is SORA from OpenAI [sora]. The videos produced from SORA are of extremely high-fidelity that makes it nearly indistinguishable from real-life footage. Thus, SORA brings in new challenges in tackling misinformation, identifying deepfakes, and distinguishing real from fake videos especially during crucial events like democratic elections. Therefore, developing automated solutions to identify AI-generated videos has become the need of the hour.
@@
Motivated by the above mentioned points, we propose a novel and challenging benchmark, VANE-Bench: Video ANomaly Evaluation Benchmark, to evaluate various closed-source, and open-source Video-LMMs on their ability to detect anomalies in the videos. Our VANE-Bench consists of both real-world video anomalies from diverse surveillance footage capturing unusual pedestrian behaviour, criminal activities, and unusual events, as well as subtle and challenging anomalies and inconsistencies present in various AI-generated videos (See Fig 1). These AI-generated videos, especially from SOTA models like SORA, have subtle and hard to detect anomalies, which makes this a challenging task even for many humans. However, automatically detecting and identifying the anomalies in these synthetic video clips serves as an important step towards identifying AI-generated videos in the wild. We reformulate the problem statement of VAD into a visual question-answering (VQA) task to facilitate easier evaluation of LMMs. However, despite evaluating over nine recent Video-LMMs on VANE-Bench, we find that most current LMMs still struggle on this benchmark (see Fig.2), making VANE-Bench a challenging and a useful benchmark for tracking the progress of Video-LMMs for the foreseeable future.
@@
    We open-source our code, and describe the data construction process of VANE-Bench along with making our data publicly available.
@@
Figure 2: Left: Performance of Video-LMMs on five anomaly categories of SORA dataset. Right: Overall performance of Video-LMMs averaged across all the benchmark datasets including AI-generated and real-world anomaly datasets.
@@ 2 Related Work
Video-LMMs: LMMs integrate linguistic and visual data to process videos, leveraging LLMs like Llama [llama3] and connecting them with modality-specific encoders via interfaces like Q-former [zhang2024mmllms, dai2023instructblip, yin2024survey]. Notable open-source Video-LMMs include VideoChat [li2023videochat], which uses a chat-centric system, and VideoChatGPT [maaz2023video], which combines a visual encoder with an LLM for detailed video conversations. Video-LLaMA [zhang2023video] integrates audio and visual signals using Q-formers, while LLaMA-VID [li2023llama] represents each frame with context and content tokens for efficient processing. Despite these advancements, our work shows current LMMs perform poorly on VANE-Bench, highlighting the need for stronger models in anomaly detection.
Video-LMMs Benchmarking: Benchmarks like SEED-Bench [li2023seedbench] and MV-Bench [li2024mvbench] assess general comprehension through multiple-choice questions but lack focus on anomaly detection in AI-generated videos. CVRR-ES [Khattak2024cvrres] evaluates real-world scenarios with open-ended questions but doesn’t address AI-generated inconsistencies. VANE-Bench specifically evaluates VAD in both real-world and AI-generated videos, providing a targeted benchmark for this task. While Perception Test [pătrăucean2023perception] focuses on lower-level perception in real-world videos, VANE-Bench targets subtle anomalies in AI-generated content, making it essential for assessing Video-LMM robustness.
Video Anomaly Detection: Traditional VAD methods typically rely on hand-crafted features and statistical models to identify deviations from normality. CUVA [du2024uncovering] is a comprehensive benchmark which focuses on the causation of video anomalies. A survey on generalized VAD [liu2024generalized] categorizes various methodologies and highlights benchmark limitations. These methods often fail with complex AI-generated videos. VANE-Bench addresses this by focusing on VAD in such videos, complementing existing benchmarks and targeting subtle inconsistencies in high-fidelity AI-generated content.
@@ 3 Dataset & Benchmark
Recent advancements in multi-modal Large Language Models (LLMs) have enabled these models to process text, image, and video data, presenting new opportunities and challenges in Video Anomaly Detection (VAD) [vad-survey]. Motivated by this progress, we aim to benchmark the capabilities of these multi-modal models (LMMs) on VAD.
To address VAD, we propose VANE-Bench: Video ANomaly Evaluation Benchmark for Conversational LMMs, comprising 325 video clips and 559 challenging ground-truth question-answer (QA) pairs. We have adapted the VAD problem into a Multiple-Choice Video Question Answering (MC-Video QA) [tapaswi2016movieqa, lei2019tvqa, yu2019activitynetqa] task to facilitate the evaluation of LMMs, allowing for a more granular assessment of their video content understanding.
@@ 3.1 Overview
VANE-Bench consists of 325 video clips spanning real-world and synthetic video anomalies. We adapted standard VAD surveillance datasets such as CUHK Avenue [cuhk-avenue], UCF-Crime [ucf-crime], and UCSD Pedestrian [ucsd-ped] to our MC-Video QA problem. Additionally, we included 197 video clips from various open-source and state-of-the-art closed-source text-to-video diffusion models [sora, open-sora, gen2-runway, wang2023modelscope, wang2023videolcm].
@@ 3.2 Categories
Figure 3: VANE-Bench dataset statistics: Left and Middle: Composition and type of anomalies present in AI-generated and real-world video samples. Right: Number of samples and QA pairs present in each type of video dataset.
@@
  • Real-World anomalies: The videos with these anomalies are sourced from several established real-world anomaly datasets, encompassing diverse anomaly types. The distribution of these anomalies is depicted in Fig. 3 (left). Fig. 3 (right) provides the total number of anomaly clips along with corresponding QA pairs for each dataset in this category. Detailed descriptions of each dataset within this category follow below.
@@
      CUHK Avenue [cuhk-avenue]: This category consists of 11 video clips with 33 associated question-answer (QA) pairs. The clips capture anomalous events in a campus environment, which shows individuals commuting in a university campus, and walking in and out of buildings. Anomaly types. The anomalies include unusual pedestrian behavior like randomly throwing bags and papers, or performing weird actions or dance moves.
@@
      UCF-Crime [ucf-crime]: Comprising 95 video clips with 95 QA pairs, this category includes real-world surveillance footage. Anomaly types. The videos depict various criminal activities, such as arrest, assault, burglary, robbery, stealing, and vandalism.
@@
      UCSD-Ped1 [ucsd-ped]: This category contains 10 video clips with 30 QA pairs. The videos focus on pedestrian walkways. The Ped1 dataset is captured by a camera facing perpendicular to the road. Anomaly types. The anomalous events are due to the presence of non pedestrian entities (i.e. bikers, skaters, small carts, and wheelchairs) in the walkways.
@@
      UCSD-Ped2 [ucsd-ped]: Similar to UCSD-Ped1, this category includes 12 video clips with 36 QA pairs. In contrast with Ped1, the Ped2 dataset uses camera which is parallel to the road. Anomaly types. Abnormal events are due to non pedestrian entities in the walkways including bikers, skaters, small carts, and people walking across a walkway.
  • AI-Generated anomalies: The videos with these anomalies are obtained from various closed-source, and open-source text-to-video diffusion models. The anomalies in these clips are usually subtle, and hard to detect, which makes our VANE-Bench benchmark challenging. General anomaly types: The anomalies include sudden appearance of objects, unnatural transformation of solid physical objects, disappearance of objects, objects passing through other solids, and unnatural appearance of objects (i.e. distorted and deformed facial features, or other unnatural appearance like presence of extra fingers). The distribution of these anomalies in the dataset is shown in Fig. 3 (middle) and statistics about the number of clips and corresponding QA pairs is presented in Fig. 3 (right). Below we describe type of video samples in this category.
@@
      SORA [sora]: This category consists of 46 video clips with 138 QA pairs. The video clips are generated using SORA, a state-of-the-art AI text-to-video model. Due to the high quality, and almost realistic looking videos generated by SORA, it becomes quite difficult to accurately identify the inconsistencies or anomalies present in the videos.
@@
      OpenSora [open-sora]: With 50 video clips and 50 QA pairs, this category features AI-generated videos from the open-source version of SORA.
@@
      Runway Gen2 [gen2-runway]: This category includes 25 video clips with 25 QA pairs, created using a commercial text-to-video AI model.
@@
      ModelScopeT2V [wang2023modelscope]: This category comprises of 24 video clips with 48 QA pairs, leveraging the video diffusion model trained by [wang2023modelscope] to produce videos from text captions. The videos were generated with 50 diffusion steps with 16 fps.
@@
      VideoLCM [wang2023videolcm]: This category features 52 video clips with 104 QA pairs, generated using latent consistency models [wang2023videolcm] designed to create videos with high variability and with less latency. We used 20 diffuson steps to generate the videos with 16 fps. The videos were further post-processed by an LCM model trained on higher resolution videos to obtain better quality videos.
@@ 3.3 Constructing VANE-Bench
Fig. 4 describes the construction process of the VANE-Bench dataset. Since the synthetic AI generated videos from state-of-the-art models like SORA [sora] have subtle, and hard to detect inconsistencies, we require high quality captions describing all of the specific inconsistencies present in the given video. Our pipeline first annotates the anomalies using the frame annotation module (FAM). The caption generating module (CGM) then utilizes these annotations to produce captions, followed by the question-answer generation module (QAGM) creating QA pairs based on the annotated frames and captions. Annotating the clips before caption generation is crucial for focusing the model on the specific anomaly regions in the video [shtedritski2023does, zhang2023gpt4roi, yang2023set]. Without annotations, the CGM often fails to reference the anomalies in the captions, as demonstrated in Sec. C of supplementary material. We briefly describe all the three stages involved in the semi-automatic dataset construction pipeline below.

### v2

# VANE-Bench: Video Anomaly Evaluation Benchmark for Conversational LMMs

@@ Abstract
The recent advancements in Large Language Models (LLMs) have greatly influenced the development of Large Multi-modal Video Models (Video-LMMs), significantly enhancing our ability to interpret and analyze video data. Despite their impressive capabilities, current Video-LMMs have not been evaluated for anomaly detection tasks, which is critical to their deployment in practical scenarios e.g., towards identifying deepfakes, manipulated video content, traffic accidents and crimes. In this paper, we introduce VANE-Bench, a benchmark designed to assess the proficiency of Video-LMMs in detecting and localizing anomalies and inconsistencies in videos. Our dataset comprises an array of videos synthetically generated using existing state-of-the-art text-to-video generation models, encompassing a variety of subtle anomalies and inconsistencies grouped into five categories: unnatural transformations, unnatural appearance, pass-through, disappearance and sudden appearance. Additionally, our benchmark features real-world samples from existing anomaly detection datasets, focusing on crime-related irregularities, atypical pedestrian behavior, and unusual events. The task is structured as a visual question-answering challenge to gauge the models’ ability to accurately detect and localize the anomalies within the videos. We evaluate nine existing Video-LMMs, both open and closed sources, on this benchmarking task and find that most of the models encounter difficulties in effectively identifying the subtle anomalies. In conclusion, our research offers significant insights into the current capabilities of Video-LMMs in the realm of anomaly detection, highlighting the importance of our work in evaluating and improving these models for real-world applications. Our code and data is publicly available at https://github.com/rohit901/VANE-Bench.
VANE-Bench: Video Anomaly Evaluation Benchmark
for Conversational LMMs
Hanan Gani*1  Rohit Bharadwaj*2  Muzammal Naseer3 Fahad Shahbaz Khan1,4Salman Khan1,5 1Mohamed Bin Zayed University of Artificial Intelligence, 2University of Edinburgh, 3Department of Computer Science, Khalifa University, 4Linköping University, 5Australian National University Correspondence: hanan.ghani@mbzuai.ac.ae, rohit.bharadwaj@ed.ac.uk
**footnotetext: Equal contribution
Figure 1: Samples showing the AI-Generated video category of VANE-Bench. We collect these synthetic videos from SORA Brooks et al. (2024), Open-Sora HPCAI Tech (2024), Runway Gen2 Runway Research (2024), ModelScopeT2V Wang et al. (2023a), and VideoLCM Wang et al. (2023b). The correct option in each question is highlighted in bold. Note that many of these anomalies are extremely subtle and difficult for humans to detect since the changes happen in rapid succession, with the entire video played in under a second. Anomalies are identified with red bounding boxes for clarity. Note that our actual dataset does not contain bounding box overlays.
@@ 1 Introduction
Large Language Models (LLMs) like ChatGPT have ushered in a new era of real-world AI applications in varied and diverse sectors like manufacturing, legal services, space exploration, transportation, retail, healthcare, education, and technology Abdullah et al. (2022); Marr (2023). Further, the current trend in the development of these LLMs has been to introduce multi-modal capabilities like vision and audio to these models along with text et al. (2024); OpenAI (2024). This motivates us to ask the question whether the current Large Multi-modal Models (LMMs) are capable and accurate in tackling the problem statement of Video Anomaly Detection (VAD) which has immense practical applications in factories, autonomous driving, crime warning, and traffic management Liu et al. (2024a).
@@
Further, we have recently observed superior visual quality of various AI-generated videos due to the improvements in the underlying algorithms, which are based on diffusion models, and transformers Brooks et al. (2024); HPCAI Tech (2024); Peebles and Xie (2023). The current state-of-the-art (SOTA) AI text-to-video model is SORA from OpenAI Brooks et al. (2024). The videos produced by SORA are of extremely high fidelity, which makes them nearly indistinguishable from real-life footage. Thus, SORA brings new challenges in tackling misinformation, identifying deepfakes, and distinguishing real from fake videos, especially during crucial events like democratic elections. Therefore, developing automated solutions to identify AI-generated videos has become the need of the hour.
@@
Motivated by the above-mentioned points, we propose a novel and challenging benchmark, VANE-Bench: Video ANomaly Evaluation Benchmark, to evaluate various closed-source and open-source Video-LMMs on their ability to detect anomalies in the videos. Our VANE-Bench consists of both real-world video anomalies from diverse surveillance footage capturing unusual pedestrian behaviour, criminal activities, and unusual events, as well as subtle and challenging anomalies and inconsistencies present in various AI-generated videos (See Fig 1). These AI-generated videos, especially from SOTA models like SORA, have subtle and hard to detect anomalies, which makes this a challenging task even for many humans. However, automatically detecting and identifying the anomalies in these synthetic video clips serves as an important step towards identifying AI-generated videos in the wild. We reformulate the problem statement of VAD into a visual question-answering (VQA) task to facilitate easier evaluation of LMMs. However, despite evaluating over nine recent Video-LMMs on VANE-Bench, we find that most current LMMs still struggle on this benchmark (see Fig.2), making VANE-Bench a challenging and a useful benchmark for tracking the progress of Video-LMMs for the foreseeable future.
@@
    We will open-source our code, and describe the data construction process of VANE-Bench along with making our data publicly available.
@@
Figure 2: Left: Performance of Video-LMMs on five anomaly categories of SORA dataset. Right: Overall performance of Video-LMMs averaged across all the benchmark datasets, including AI-generated and real-world anomaly datasets.
@@ 2 Related Work
Video-LMMs: LMMs integrate linguistic and visual data to process videos, leveraging LLMs like Llama Meta (2024) and connecting them with modality-specific encoders via interfaces like Q-former Zhang et al. (2024); Dai et al. (2023); Yin et al. (2024). Notable open-source Video-LMMs include VideoChat Li et al. (2023b), which uses a chat-centric system, and VideoChatGPT Maaz et al. (2023), which combines a visual encoder with an LLM for detailed video conversations. Video-LLaMA Zhang et al. (2023a) integrates audio and visual signals using Q-formers, while LLaMA-VID Li et al. (2023c) represents each frame with context and content tokens for efficient processing. Despite these advancements, our work shows current LMMs perform poorly on VANE-Bench, highlighting the need for stronger models in anomaly detection.
Video-LMMs Benchmarking: Benchmarks like SEED-Bench Li et al. (2023a) and MV-Bench Li et al. (2024b) assess general comprehension through multiple-choice questions but lack focus on anomaly detection in AI-generated videos. CVRR-ES khattak et al. (2024) evaluates real-world scenarios with open-ended questions but doesn’t address AI-generated inconsistencies. VANE-Bench specifically evaluates VAD in both real-world and AI-generated videos, providing a targeted benchmark for this task. While Perception Test Pătrăucean et al. (2023) focuses on lower-level perception in real-world videos, VANE-Bench targets subtle anomalies in AI-generated content, making it essential for assessing Video-LMM robustness.
Video Anomaly Detection: Traditional VAD methods typically rely on hand-crafted features and statistical models to identify deviations from normality. CUVA Du et al. (2024) is a comprehensive benchmark that focuses on the causation of video anomalies. A survey on generalized VAD Liu et al. (2024b) categorizes various methodologies and highlights benchmark limitations. These methods often fail with complex AI-generated videos. VANE-Bench addresses this by focusing on VAD in such videos, complementing existing benchmarks and targeting subtle inconsistencies in high-fidelity AI-generated content.
@@ 3 Dataset & Benchmark
Recent advancements in multi-modal Large Language Models (LLMs) have enabled these models to process text, image, and video data, presenting new opportunities and challenges in Video Anomaly Detection (VAD) Liu et al. (2024a). Motivated by this progress, we aim to benchmark the capabilities of these multi-modal models (LMMs) on VAD.
To address VAD, we propose VANE-Bench: Video ANomaly Evaluation Benchmark for Conversational LMMs, comprising 325 video clips and 559 challenging ground-truth question-answer (QA) pairs. We have adapted the VAD problem into a Multiple-Choice Video Question Answering (MC-Video QA) Tapaswi et al. (2016); Lei et al. (2019); Yu et al. (2019) task to facilitate the evaluation of LMMs, allowing for a more granular assessment of their video content understanding.
@@ 3.1 Overview
VANE-Bench consists of 325 video clips spanning real-world and synthetic video anomalies. We adapted standard VAD surveillance datasets such as CUHK Avenue Lu et al. (2013), UCF-Crime Sultani et al. (2018), and UCSD Pedestrian Li et al. (2014) to our MC-Video QA problem. Additionally, we included 197 video clips from various open-source and state-of-the-art closed-source text-to-video diffusion models Brooks et al. (2024); HPCAI Tech (2024); Runway Research (2024); Wang et al. (2023a, b).
@@ 3.2 Categories
Figure 3: VANE-Bench dataset statistics: Left and Middle: Composition and type of anomalies present in AI-generated and real-world videos. Right: Number of samples and QA pairs present in each type of video dataset.
@@
Real-World anomalies: The videos with these anomalies are sourced from several established real-world anomaly datasets, encompassing diverse anomaly types. The distribution of these anomalies is depicted in Fig. 3 (middle). Fig. 3 (right) provides the total number of anomaly clips along with corresponding QA pairs for each dataset in this category. Detailed descriptions of each dataset within this category follow below.
@@
    CUHK Avenue Lu et al. (2013): This category consists of 11 video clips with 33 associated question-answer (QA) pairs. The clips capture anomalous events in a campus environment, which shows individuals commuting in a university campus, and walking in and out of buildings. Anomaly types. The anomalies include unusual pedestrian behavior like randomly throwing bags and papers or performing weird actions or dance moves.
@@
    UCF-Crime Sultani et al. (2018): Comprising 95 video clips with 95 QA pairs, this category includes real-world surveillance footage. Anomaly types. The videos depict various criminal activities, such as arrest, assault, burglary, robbery, stealing, and vandalism.
@@
    UCSD-Ped1 Li et al. (2014): This category contains 10 video clips with 30 QA pairs. The videos focus on pedestrian walkways. The Ped1 dataset is captured by a camera facing perpendicular to the road. Anomaly types. The anomalous events are due to the presence of non pedestrian entities (i.e. bikers, skaters, small carts, and wheelchairs) in the walkways.
@@
    UCSD-Ped2 Li et al. (2014): Similar to UCSD-Ped1, this category includes 12 video clips with 36 QA pairs. In contrast with Ped1, the Ped2 dataset uses camera which is parallel to the road. Anomaly types. Abnormal events are due to non pedestrian entities in the walkways including bikers, skaters, small carts, and people walking across a walkway.
AI-Generated anomalies: The videos with these anomalies are obtained from various closed-source, and open-source text-to-video diffusion models. The anomalies in these clips are usually subtle, and hard to detect, which makes our VANE-Bench benchmark challenging. General anomaly types: The anomalies include the sudden appearance of objects, the unnatural transformation of solid physical objects, the disappearance of objects, objects passing through other solids, and unnatural appearance of objects (i.e., distorted and deformed facial features, or other unnatural appearance like presence of extra fingers). The distribution of these anomalies in the dataset is shown in Fig. 3 (left), and statistics about the number of clips and corresponding QA pairs are presented in Fig. 3 (right). Below, we describe the type of video samples in this category.
@@
    SORA Brooks et al. (2024): This category consists of 46 video clips with 138 QA pairs. The video clips are generated using SORA, a state-of-the-art AI text-to-video model. Due to the high quality and almost realistic-looking videos generated by SORA, it becomes quite difficult to accurately identify the inconsistencies or anomalies present in the videos.
@@
    OpenSora HPCAI Tech (2024): With 50 video clips and 50 QA pairs, this category features AI-generated videos from the open-source version of SORA.
@@
    Runway Gen2 Runway Research (2024): This category includes 25 video clips with 25 QA pairs created using a commercial text-to-video AI model.
@@
    ModelScopeT2V Wang et al. (2023a): This category comprises of 24 video clips with 48 QA pairs, leveraging the video diffusion model trained by Wang et al. (2023a) to produce videos from text captions. The videos were generated with 50 diffusion steps with 16 fps.
@@
    VideoLCM Wang et al. (2023b): This category features 52 video clips with 104 QA pairs, generated using latent consistency models Wang et al. (2023b) designed to create videos with high variability and with less latency. We used 20 diffuson steps to generate the videos with 16 fps. The videos were further post-processed by an LCM model trained on higher resolution videos to obtain better quality videos.
@@ 3.3 Constructing VANE-Bench
Fig. 4 describes the construction process of the VANE-Bench dataset. Since the synthetic AI-generated videos from state-of-the-art models like SORA Brooks et al. (2024) have subtle and hard-to-detect inconsistencies, we require high-quality captions describing all of the specific inconsistencies present in the given video. Our pipeline first annotates the anomalies using the frame annotation module (FAM). The caption-generating module (CGM) then utilizes these annotations to produce captions, followed by the question-answer generation module (QAGM), creating QA pairs based on the annotated frames and captions. Annotating the clips before caption generation is crucial for focusing the model on the specific anomaly regions in the video Shtedritski et al. (2023); Zhang et al. (2023b); Yang et al. (2023). Without annotations, the CGM often fails to reference the anomalies in the captions, as demonstrated in Sec. C of supplementary material. We briefly describe all the three stages involved in the semi-automatic dataset construction pipeline below.